Cost Behavior, High-Low Method, Pricing Decision

St. Teresa's Medical Center (STMC) offers a number of specialized medical services, including neuroscience, cardiology, and oncology. STMC's strong reputation for quality medical care allowed it to branch out into other services. It is now ready to expand its orthopedic services and has just added a free-standing orthopedic clinic offering a full range of outpatient, surgical, and physical therapy services. The cost of the orthopedic facility is depreciated on a straight-line basis. All equipment within the facility is leased.

Since the clinic had no experience with in-patient orthopedic services (for patients recovering from hip and knee replacements, for example), it decided to operate the orthopedic center for two months before determining how much to charge per patient day on an ongoing basis. As a temporary measure, the clinic adopted a patient-day charge of $190, an amount equal to the fees charged by a hospital specializing in orthopedic care in a nearby city.

This initial per-day charge was quoted to patients entering the orthopedic center during the first two months with assurances that if the actual operating costs of the new center justified it, the charge could be less. In no case would the charges be more. A temporary policy of billing after 60 days was adopted so that any adjustments could be made.

The orthopedic center opened on January 1. During January, the center had 4,200 patient days of activity. During February, the activity was 4,600 patient days. Costs for these two levels of activity output are as follows:

4,200 Patient Days 4,600 Patient Days
Salaries, nurses $ 59,500 $ 59,500
Aides 30,700 30,700
Pharmacy 246,300 267,900
Laboratory 99,300 106,500
Depreciation 28,000 28,000
Laundry 189,840 207,920
Administration 23,000 23,000
Lease (equipment) 37,900 37,900

1. Use the high-low method to separate the mixed costs into fixed and variable.

Laboratory: Pharmacy:
Variable $ per patient day $ per patient day
Fixed $ $

2.  The administrator of the orthopedic center estimated that the center will average 4,400 patient days per month. If the center is to be operated as a nonprofit organization, determine the amount it will need to charge per patient day? Round your interim calculations and final answers to the nearest cent.

$ Charge per patient day

How much of this charge is variable?

$ Variable charge per patient day

How much of the charge per patient day is fixed?

$ Fixed charge per patient day

3. Suppose the orthopedic center averages 4,800 patient days per month. How much would need to be charged per patient day for the center to cover its costs? Round your answer to the nearest cent.

$ per patient day

The main reason why the charge per patient day decreased as the activity output increased is because:
Answer #1

Ans.1. Laboratory- Increase in Patient Days= 4600 - 4200 = 400 Increase in Cost= 106,500 - 99,300 = $7,200 Therefore, Variable Cost per Patient Day = ($7,200 / 400) = $18/patient day and, Fixed Cost = $99,300 - ($18/patient day x 4,200) = $99,300 - $75,600 = $23,700

Pharmacy- Increase in Patient Days= 4600 - 4200 = 400 Increase in Cost= 267,900 - 246,300 = $21,600 Therefore, Variable Cost per Patient Day = ($21,600 / 400) = $54/patient day and, Fixed Cost = $246,300 - ($54/patient day x 4,200) = $246,300 - $226,800 = $19,500

Laboratory Pharmacy
Variable $18 per patient day $54 per patient day
Fixed $23,700 $19,500

Ans.2.

Cost Estimation for 4,400 Patient Days
Particulars Variable Fixed Total
Salaries, nurse 59,500 59,500
Add: Aides 30,700 30,700
Add: Laboratory 79,200 23,700 102,900
Add: Pharmacy 237,600 19,500 257,100
Add: Depreciation 28,000 28,000
Add: Laundry 198,880 198,880
Add: Admin 23,000 23,000
Add: Lease (Equipment) 37,900 37,900
Total 737,980

Charge per patient day = $737,980 / 4,400 = $167.72 Variable Cost % = 69.88% Variable Cost per patient day = $117.20 Fixed Cost % = 30.12% Fixed Cost per patient day = $50.52

Ans.3.

Cost Estimation for 4,800 Patient Days
Particulars Variable Fixed Total
Salaries, nurse 59,500 59,500
Add: Aides 30,700 30,700
Add: Laboratory 86,400 23,700 110,100
Add: Pharmacy 259,200 19,500 278,700
Add: Depreciation 28,000 28,000
Add: Laundry 216.960 216.960
Add: Admin 23,000 23,000
Add: Lease (Equipment) 37,900 37,900
Total 784,860

Charge per patient day = $784,860 / 4,800 = $163.51

The main reason why the charge per patient day decreased as the activity output increased is because of the variable cost per patient day remaining same and the fixed cost per patient day reducing due to increase in output.